OpenStudy (anonymous):

The formula for the volume of a rectangular box is V = LWH. If V = 200, L= 12.5, and W = 8, find the height of the box, to the nearest whole unit.

OpenStudy (ash2326):

V=LWH 200=12.5*8*H 200=100H H=2

OpenStudy (anonymous):

Plug the values into V = LWH 200 = 12.5*8*H

OpenStudy (anonymous):

200 = 12.5 * 8 * H 200 = 100H 200/100 = H 2 = H H = 2